Is the sequence 〖{a〗_n} a solution of the recurrence relation a_n=8a_(n-1)-16a_(n-2)if a_n=n^4 a_n=0.
1
Expert's answer
2020-06-01T17:11:47-0400

Solution.

"a_n=8a_{n-1}-16a_{n-2};"

a) "a_n=0?"

Replace "n" in "a_n=0" by "n-1"

"a_{n-1}=0";

Replace "n" in "a_n=0" by "n-2"

"a_{n-2}=0;"

"8a_{n-1}-16a_{n-2}=8\sdot0-16\sdot 0=0=a_n"; Yes

b)"a_n=n^4?"

Replase "n" in "a_n=n^4" by "n-1"

"a_{n-1}=(n-1)^4";

Replase "n" in "a_n=n^4" by "n-2"

"a_{n-2}=(n-2)^4"

"8(n-1)^4-16(n-2)^4=8(n^4-4n^3+6n^2-4n+1)-16(n^4-8n^3+24n^2+-32n+16)=8n^4-32n^3+48n^2-32n+8-16n^4+128n^3-384n^2+512n-256=-8n^4+96n^3-336n^2+480n-248{=}\mathllap{/\,}n^4;"

No

Answer: a)Yes;

b)No.